OpenStudy (moneybaby980):

Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in Choose one…. September October by Choose one…. $30 $75 $190

OpenStudy (anonymous):

Are you told the selling price of each flower in Part A maybe?

OpenStudy (moneybaby980):

yes

OpenStudy (anonymous):

So what is it?

OpenStudy (moneybaby980):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $

OpenStudy (moneybaby980):

$1.25

OpenStudy (anonymous):

Well that's in October heh, now I see that there is a different price for september

OpenStudy (anonymous):

Ok, so now let's see, what is 40% of $1.50?

OpenStudy (moneybaby980):

2.5

OpenStudy (anonymous):

40% of something cannot be bigger than that thing

OpenStudy (moneybaby980):

0.6

OpenStudy (anonymous):

First let's calculae 1% of 1.50, which will be \( 1\% = \frac{1.50}{100} = 0.015\) Now let's calculate 40, which will be 40 times 1% so \(40\% = 0.015 \times 40 = 0.6 \)

OpenStudy (anonymous):

right, so they earned 0.6 of each flower sold in october. let's see what is 40% of 1.25 as well

OpenStudy (moneybaby980):

0.5

OpenStudy (anonymous):

Exactly. so they earned $0.5 for every flower in september. Now we are told they sold 900 flowers in september so it becomes \(0.5 \times 900 = 450\) In october we calculated that they earned $0.6 for each flower, we are told they wold 700. so how much did they earn?

OpenStudy (moneybaby980):

420

OpenStudy (anonymous):

Right, so which they earned more? what's the difference?

OpenStudy (moneybaby980):

September by $30

OpenStudy (anonymous):

Right =)